# Black hole surrounded by a magnetic vortex in $f(R)$ gravity

### v1

@@ Abstract
Abstract: Modified Gravity Theories (MGTs) are extensions of General Relativity (GR) in its standard formulation. In light of this premise, we employ Heisenberg’s non-perturbative approach to promote the three-dimensional Einstein-Hilbert theory to a modified $f(R)$ gravity. Within this context, we investigate a cosmological system composed of a black hole (BH) surrounded by Maxwell-Higgs vortices, forming the BH-vortex system. In the case of linear $f(R)$ gravity derived from quantum metric fluctuations, one shows the existence of a three-dimensional ring-like BH-vortex system with quantized magnetic flux. Within this system, one notes the BH at $r=0$ and its event horizon at $r=r_{0}$, while the magnetic vortices are at $r\in(r_{0},\infty)$. A remarkable result is the constancy of the Bekenstein-Hawking temperature ($T_{H}$), regardless of MGTs and vortex parameters. This invariance of $T_{H}$ suggests that the BH-vortex system reaches thermodynamic stability. Unlike the standard theory of Maxwell-Higgs vortices in flat spacetime, in $f(R)$ gravity, the vortices suffer the influence of the BH’s event horizon. This interaction induces perturbations in the magnetic vortex profile, forming cosmological ring-like magnetic structures.
Keywords: Modified Gravity Theory. Metric Fluctuation. Black Holes. Magnetic Vortices.
@@ I Introduction
Recently, we noted a growing interest in Modified Gravity Theories (MGTs). Generally speaking, these theories arise from the premise that the universe is suffering an accelerated expansion Tonry , Knop , which can be explained through corrections to the equations of motion in General Relativity (GR) GOlmo . Initially, one explanation for this universal acceleration is the hypothesis that additional contributions from the scalar curvature may exist in the standard formulation of GR, leading to MGTs, e.g., $f(R)$ theory SCapozziello , SMCarroll1 , Vollick , SMCarroll2 . Specifically, one of the motivations for $f(R)$ gravity is the study of possible new gravitational effects related to late-time cosmic expansion, see Refs. SCapozziello3 , Clifton , Eingorn , Saffari , Capozziello4 , Sotirou . These remarks have provided extensive studies on MGTs as promising frameworks for addressing emerging challenges to standard GR Almeida2 . Among these theories are $f(R)$ Felice , SNojiriAndOdintsov , SCapozziello4 and its variants as $f(R,T)$ gravity THarkoX , GoncalvesX , Rosa1X , Rosa2X , models involving couplings between curvature and matter THarko0 , and proposals based on modified geometries Lavinia , Cai . Thus, this jungle of MGTs has found significant applications in cosmological and astrophysical contexts Saito1 , Astashenok , Davis .
In the framework of MGTs, a fundamental question arises: which of these theories is most appropriate for describing systems composed of gravitating compact objects? Traditionally, one chooses the MGT by basing on constraints in weak-field regimes derived from classical tests of GR applied to models similar to the solar system Chiba1 , Erickcek , Chiba2 , NojiriX , CapozzielloX . Thus, we propose an alternative approach that departs from this conventional perspective. Our goal is to determine the most suitable gravity theory for studying a system composed of a black hole and magnetic vortex (BH-vortex). Towards this purpose, one adopts an approach based on quantum fluctuations of the metric within the Einstein-Hilbert theory to identify the MGT capable of adequately describing this system. Summarizing the quantum fluctuation approach considers a metric that incorporates classical ($g^{\mu\nu}$) and quantum ($\delta\hat{g}^{\mu\nu}$) contributions Dzhu1 , Dzhu2 , Dzhu3 . This approach offers significant advantages as it naturally introduces modifications to Einstein’s gravity, allowing, among other aspects, the emergence of non-minimal couplings between gravity and matter. By adopting this methodology, we assume that the expected value of the metric satisfies certain physical conditions, such as the non-vanishing of the average quantum contribution and the ability to express the metric $g^{\mu\nu}$ in terms of a second-rank symmetric tensor Dzhu1 , Dzhu2 , Dzhu3 . These assumptions lead to a theoretical scenario in which the modified gravity of the $f(R)$ emerges as a fundamental structure for describing BH-vortex systems. Grounded in this perspective, our objective in this work is to analyze a self-gravitating three-dimensional BH-vortex system in $f(R)$ gravity. Fig. 1 illustrates the system under consideration.
Figure 1: Illustration of the BH-vortex system.
@@
In several contexts, one adopts widely the MGTs Zhdanov , Nojiri00 , Giacomozzi , Nashed ; however, one knows little about the self-gravitating topological structures [^1] Topological structures are nonlinear field configurations with stability due to topological properties. These objects arise particularly in field theory and particle physics. Generally speaking, one classifies these structures as kinks, vortices, and monopoles Vachaspati , Rajaraman , Manton , Vilenkin .. A significant class of topological structures are the topological vortices, which emerge from a complex scalar field theory coupled to an Abelian gauge field Vachaspati , Rajaraman , Manton , Vilenkin . Initially, Higgs used a massive scalar field in the mechanism that would later bear his name Higgs . Subsequently, the investigation of models coupled to an Abelian gauge field was introduced by Nielsen and Olesen Nielsen , adopting a flat spacetime. Generally speaking, vortex models preserve the Higgs mechanism and gauge invariance and are particularly relevant once these models are relativistic generalizations of the phenomenological Ginzburg-Landau theory for superconductivity James . Thus, the following question naturally arises: would it be possible for a BH-vortex system to emerge in an MGT? We aim to address this question in the development of this article.
Magnetic vortices arise in condensed matter physics Abrikosov , Albert ; however, their application has expanded to describe vortex lines in cosmic strings Nielsen , Albert . Recently, vortices have gained increasing interest due to their reinterpretation within a cosmological framework, particularly in the BHs description Dvali , Lima . In this context, vortices are intrinsically related to the BH dynamic, enabling the emission or capture of magnetic matter fluxes Lima . The study of the BH-vortex system can thus provide a topological interpretation for the stability of BHs and their implications in Hawking radiation Dvali . Furthermore, one performs several studies on BHs due to their relevance in contemporary physics Frolov and experimental findings Akiyama . Thus, investigations on these objects have grown significantly, see Refs. Choudhury , Ditta , Zhu , Kehagias . In simple terms, BHs arise as solutions from the Einstein field equations Rindler . The first BH solution was announced by K. Schwarzschild Schwarzschild , describing static and spherical BH. Afterward, Oppenheimer and Snyder showed that Schwarzschild BH solution represents the final state of the collapse of a massive and spherical star Oppenheimer .
Although widely studied the non-perturbative vortex solutions, the investigation of BH-vortex cosmological systems remains largely unexplored. That is because the description of vortices in a flat spacetime presents a high level of complexity in their description. Additionally, from a gravitational perspective, Einstein’s gravity is considered trivial, i.e., outside localized sources, spacetime in the vacuum is locally flat Edery , Deser . Thus, meaningful discussions have emerged over time. Among these studies, the BTZ BH solution in an AdS3 background highlights an interesting topic in standard Einstein gravity within a three-dimensional spacetime Banados1 , Banados2 . Besides, Cardoni et al. Cardoni examined and announced the solutions for spherical scalar hair black holes in Einstein gravity. Naturally, one adopts a real scalar and complex scalar field with interaction by allowing the emergence of topological vortex in BHs within Einstein’s gravity. Taking that into account, we announce for the first time the study of a BH-vortex system (Fig. 1) in the MGTs framework.
We organized the manuscript into four distinct sections. We begin applying the Heisenberg formalism to the BH-vortex system (Sec. II), providing the most suitable MGT framework. This analysis leads to an $f(R)$ theory to study the BH-vortex system. Subsequently, in Sec. III, one explores the BH-vortex system within the $f(R)$ gravity. One divides this section into three subsections. We begin by outlining the general framework for generating the BH-vortex system; then, we investigate the BH solution and Hawking radiation in $f(R)$ gravity. Finally, we investigate the vortices surrounding the BH. In Sec. IV, one announces our main findings.
II Heisenberg’s Approach in a Gravitational Background
Within the framework of this work, we will consider a quantization formalism in a three-dimensional gravitational background. In this framework, let us treat the metric as an ordinary field, and we will employ Heisenberg’s non-perturbative approach. This approach requires promoting the quantities $\Gamma^{\rho}\,_{\mu\nu}$, $R^{\rho}\,_{\lambda\mu\nu}$, and $R_{\mu\nu}$ to operators, i.e., $\hat{\Gamma}^{\rho}\,_{\mu\nu}$, $\hat{R}^{\rho}\,_{\lambda\mu\nu}$, and $\hat{R}_{\mu\nu}$, respectively. Inevitably, this assumption allows us to reformulate Einstein’s equation as
$\hat{G}_{\mu\nu}=\hat{R}_{\mu\nu}-\frac{1}{2}g_{\mu\nu}\hat{R}=\kappa^{2}\hat{T}_{\mu\nu}.$ (1)
It is relevant to highlight that the definitions of these quantities remain consistent even when promoted to operators. Hence, the quantities $\hat{\Gamma}^{\rho}\,_{\mu\nu}$, $\hat{R}^{\rho}\,_{\lambda\mu\nu}$, and $\hat{R}_{\mu\nu}$ are defined as natural extensions of their classical counterparts to the quantum domain, i.e.,
$\hat{R}_{\mu\nu}=\hat{R}^{\rho}\,_{\mu\rho\nu},$ (2)
$\hat{R}^{\rho}\,_{\sigma\mu\nu}=\frac{\partial\hat{\Gamma}^{\rho}\,_{\sigma\nu}}{\partial x^{\mu}}-\frac{\hat{\Gamma}^{\rho}\,_{\sigma\mu}}{\partial x^{\nu}}+\hat{\Gamma}^{\rho}\,_{\tau\mu}\hat{\Gamma}^{\tau}\,_{\sigma\nu}-\hat{\Gamma}^{\rho}\,_{\tau\nu}\hat{\Gamma}^{\tau}\,_{\sigma\mu},$ (3)
$\hat{\Gamma}^{\rho}\,_{\mu\nu}=\frac{1}{2}\hat{g}^{\rho\sigma}\left(\frac{\partial\hat{g}_{\mu\sigma}}{\partial x^{\nu}}+\frac{\partial\hat{g}_{\nu\sigma}}{\partial x^{\mu}}-\frac{\partial\hat{g}_{\mu\nu}}{\partial x^{\sigma}}\right).$ (4)
Adopting Heisenberg’s approach, one arrives at an infinite set of equations governing all Green’s functions. We express these equations as
$\begin{matrix}\langle\mathcal{Q}|\hat{g}(x_{1})\cdot(\hat{G}_{\mu\nu}-\kappa^{2}\hat{T}_{\mu\nu})|\mathcal{Q}\rangle=0;\\ \langle\mathcal{Q}|\hat{g}(x_{1})\cdot\hat{g}(x_{2})\cdot(\hat{G}_{\mu\nu}-\kappa^{2}\hat{T}_{\mu\nu})|\mathcal{Q}\rangle=0;\\ \vdots\\ \langle\mathcal{Q}|\hat{g}(x_{1})\cdot...\cdot\hat{g}(x_{i})\cdot(\hat{G}_{\mu\nu}-\kappa^{2}\hat{T}_{\mu\nu})|\mathcal{Q}\rangle=0,\end{matrix}$ (5)
where $|\mathcal{Q}\rangle$ is a quantum eigenstate [^2] A quantum eigenstate is a concept arising from quantum mechanics, referring to a state in which an observable has a well-defined value Sakurai , Shankar . Mathematically, one characterizes a quantum eigenstate by three essential properties: orthogonality, completeness, and Hermiticity Sakurai , Shankar ..
To analyze the system of equations (5), we decompose the metric operator $\hat{g}_{\mu\nu}$ into classical and quantum parts, assuming the premise that the expected value of the quantum part $\langle\hat{g}_{\mu\nu}\rangle$ non-trivial Dzhu1 , Dzhu2 , Dzhu3 , RYang . By doing so, we compute the expected value of the Lagrangian to first-order corrections in $\delta\hat{g}\,^{\mu\nu}$ for a three-dimensional Einstein-Hilbert-like action, viz.,
$S=\int\,d^{3}x\,\sqrt{-g}\left[-\frac{1}{16\pi}R+\mathcal{L}_{\text{matter}}\right].$ (6)
Within this structure, $\mathcal{L}_{\text{matter}}$ is the Matter Lagrangian density.
II.1 On the decomposition of the metric
Hereafter, let us assume that the quantum metric $\mathcal{G}_{\mu\nu}$ has classical and quantum contributions, i.e.,
$\mathcal{G}_{\mu\nu}=g_{\mu\nu}+\hat{g}_{\mu\nu},$ (7)
whose quantum contribution is non-trivial, namely,
$\langle\hat{g}_{\mu\nu}\rangle\neq 0.$ (8)
These assumptions lead to quantum corrections in the Einstein-Hilbert theory Dzhu1 , Dzhu2 , Dzhu3 , RYang . Considering first-order corrections, one obtains
$S(g+\hat{g})=\int\,d^{3}x\,\left[\mathcal{L}_{g}(g)+\frac{\delta\mathcal{L}_{g}}{\delta g^{\mu\nu}}\hat{g}_{\mu\nu}+\sqrt{-g}\mathcal{L}_{\text{matter}}(g)+\frac{\delta\sqrt{-g}\mathcal{L}_{\text{matter}}}{\delta g^{\mu\nu}}\hat{g}^{\mu\nu}\right],$ (9)
@@ II.1 On the decomposition of the metric
$\mathcal{L}_{g}(g)=-\frac{1}{16\pi}\sqrt{-g}R,$ (10)
i.e., $\mathcal{L}_{g}$ is the classical Einstein-Hilbert Lagrangian density.
Therefore, the Einstein-Hilbert theory (9), corrected with first-order quantum fluctuation terms of the metric, results in the effective theory
$S=\int\,d^{3}x\,\left[\mathcal{L}_{g}-\sqrt{-g}\mathcal{G}_{\mu\nu}K^{\mu\nu}+\frac{1}{2}\sqrt{-g}\,T_{\mu\nu}K^{\mu\nu}+\sqrt{-g}\mathcal{L}_{\text{matter}}\right].$ (11)
[…]
II.2 The case $K^{\mu\nu}=\langle\hat{g}^{\mu\nu}\rangle=\alpha g^{\mu\nu}$
Naturally, the simplest case that satisfies the criteria for the expectation value of the metric is $K^{\mu\nu}=\alpha g\,^{\mu\nu}$, where $||\alpha||\ll 1$ Almeida . Consequently, in light of this criteria, the minimally coupled Einstein-Hilbert action to the matter sector is corrected by quantum metric fluctuations, leading to the action
$S=\int\,d^{3}x\,\left[\mathcal{L}_{g}-\frac{1}{16\pi}\sqrt{-g}f(R)+\frac{1}{2}\sqrt{-g}\,T^{\mu}\,_{\mu}+\sqrt{-g}\mathcal{L}_{\text{matter}}\right].$ (12)
[…]
$S=\int\,d^{3}x\,\sqrt{-g}\left[-\frac{1}{16\pi}[R+f(R)]+\mathcal{L}_{\text{matter}}\right].$ (13)
Therefore, we obtain a modified $f(R)$ gravity. In this context, one notes that MGT naturally emerges as a consequence of the quantum fluctuations of the metric.
III The BH-vortex system in modified $f(R)$ gravity
From this stage onward, let us explore the BH and magnetic vortex solutions in modified $f(R)$ gravity. This study is particularly relevant and helps us understand the interaction between the magnetic vortex and the BH. Although this research branch is still relatively unexplored, this study provides us valuable insights into the physical aspects of the BH-vortex system in MGT.
III.1 The overall framework of the BH-Vortex system
MGTs are extensively studied, e.g., see Refs. Nojiri , Bohmer , Koyama , CapozielloX . Specifically, one employs the gravity $f(R)$ theories with coupling matter fields by adopting various dimensions to examine static and spherically symmetric field solutions Bazeia1 , Bazeia2 . This approach also allows us to understand how MGTs, defined as functions of the Ricci scalar ($R$), influence gravitating matter. Moreover, as previously mentioned, the Einstein-Hilbert theory corrected by quantum metric fluctuations takes the form presented in Eq. (13), which leads us to consider the Maxwell-Higgs action in $f(R)$ gravity, viz.,
$S=\int\,d^{3}x\,\sqrt{-g}\bigg{[}-\frac{1}{16\pi}[R+f(R)]+\frac{1}{2}(D_{\mu}\phi)^{\dagger}D^{\mu}\phi-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{\lambda}{4}(|\phi|^{2}-\nu^{2})^{2}\bigg{]},$ (14)
with $R$ the Ricci scalar, $\phi$ a complex scalar field, and $F_{\mu\nu}$ the electromagnetic field tensor, defined as $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$, where $A_{\mu}$ is the gauge field. The parameter $\nu$ denotes the Vacuum Expectation Value (VEV). Furthermore, the notation $D_{\mu}\phi$ is the covariant derivative, which minimally couples the matter field $\phi$ to the gauge field $A_{\mu}$. One defines the covariant derivative as
@@ III.1 The overall framework of the BH-Vortex system
To obtain static and topological structures in three-dimensional spacetime Lima , Edery , let us adopt the metric
@@
where $h(r)$ is the metric function, $r$ and $\theta$ are, respectively, the radial and angular variables [^6] For more details on this metric, see Refs. Lima , Edery . In Refs. Lima , Edery , the authors used this metric to examine the vortex solution in Einstein’s gravity..
@@
$R_{\mu\nu}(1+f_{R})-\frac{1}{2}g^{\mu\nu}[R+f(R)]+(\nabla_{\mu}\nabla_{\nu}-g_{\mu\nu}\nabla_{\beta}\nabla^{\beta})f_{R}=T_{\mu\nu},$ (22)
@@
III.2 The BH solution, Hawking radiation, and $f(R)$ gravity
To determine the profile of metric functions in gravity f(R) theory, we will focus our study on cases where the effects of the presence of vortices are minimally perceptible. In other words, let us analyze regions outside the BH-Vortice system, i.e., sectors where the effects are negligible. In this regime, one can assume $T_{\mu\mu}=0$ [Eq. (23)]. This simplification allows us to obtain using Eq. (22), three independent differential equations, corresponding to the components $tt$, $rr$, and $\theta\theta$, i.e.,
@@ III.2 The BH solution, Hawking radiation, and $f(R)$ gravity
$0\displaystyle 2rh^{\prime}(r)f^{\prime}_{R}+2h^{\prime}(r)f_{R}-rh^{\prime\prime}(r)+rf(R)=$ 0 . (27)
@@ III.2.1 The case $f(R)=\alpha R$
$R=-\frac{2h^{\prime}(r)}{r}-h^{\prime\prime}(r)\hskip 14.22636pt\text{and}\hskip 14.22636ptf(R)=-\alpha\left[\frac{2h^{\prime}(r)}{r}+h^{\prime\prime}(r)\right]$ (28)
@@
$(1+\alpha)rh^{\prime\prime}(r)=0\hskip 14.22636pt\text{with}\hskip 14.22636pt\alpha\neq 1.$ (29)
@@
$h(r)=r\mp r_{0},$ (30)
for all $\alpha\ll 1$.
Our purpose is to obtain BH-like compact objects. In pursuit of this, we consider only physically admissible solutions [Eq. (30)] from the theory, i.e., $h(r)=r-r_{0}$. Indeed, the function $h(r)=r-r_{0}$ represents a simple and linear form of a curvilinear coordinate function in three-dimensional spacetime. In this framework, one can interpret the parameter $r_{0}$ as the BH mass. Thus, this function constitutes a modification of the radial term of the Schwarzschild-like metric. By analogy with the Schwarzschild-like BH, this metric function profile implies that the temporal term assumes negative values for $r<r_{0}$, suggesting the existence of an event horizon at $r=r_{0}$. Furthermore, in the limit $r\to r_{0}$, the temporal contribution vanishes, ensuring an essential feature of an event horizon Carroll .
@@
Figure 2: Quadratic invariant of the Ricci tensor (blue curve) and the Kretschmann scalar (red curve) versus radial distance.
@@
III.2.2 On the Bekenstein-Hawking temperature

### v2

# Magnetized vortex in three-dimensional $\mathrm{f(R)}$ gravity

@@ Abstract
Abstract: Modified Gravity Theories (MGTs) are extensions of General Relativity (GR) in its standard formulation. Therefore, within this framework, we will investigate a system composed of a black hole (BH) surrounded by Maxwell-Higgs vortices, forming the BH-vortex system. In the case of linear $f(R)$ gravity is adopted showing the existence of a three-dimensional ring-like BH-vortex system with quantized magnetic flux. Within this system, one notes the BH at $r=0$ and its event horizon at $r=r_{0}$, while the magnetic vortices are at $r\in(r_{0},\infty)$. A remarkable result is the constancy of the Bekenstein-Hawking temperature ($T_{H}$), regardless of MGTs and vortex parameters. This invariance of $T_{H}$ suggests that the BH-vortex system reaches thermodynamic stability. Unlike the standard theory of Maxwell-Higgs vortices in flat spacetime, in $f(R)$ gravity, the vortices suffer the influence of the BH’s event horizon. This interaction induces perturbations in the magnetic vortex profile, forming cosmological ring-like magnetic structures.
@@ I Introduction
Recently, we noted a growing interest in Modified Gravity Theories (MGTs). Generally speaking, these theories arise from the premise that the universe is suffering an accelerated expansion Tonry , Knop , which can be explained through corrections to the equations of motion in General Relativity (GR) GOlmo . Initially, one explanation for this universal acceleration is the hypothesis that additional contributions from the scalar curvature may exist in the standard formulation of GR, leading to MGTs, e.g., $f(R)$ theory SCapozziello , SMCarroll1 . Specifically, one of the motivations for $f(R)$ gravity is the study of possible new gravitational effects related to late-time cosmic expansion, see Refs. SDOdin1 , SDOdin2 , SDOdin3 . These remarks have provided extensive studies on MGTs as promising frameworks for addressing emerging challenges to standard GR Almeida2 . Among these theories are $f(R)$ Felice , SCapozziello4 and its variants as $f(R,T)$ gravity THarkoX , models involving couplings between curvature and matter THarko0 , and proposals based on modified geometries Lavinia , Cai . Thus, this jungle of MGTs has found significant applications in cosmological and astrophysical contexts Saito1 .
@@
In several contexts, one adopts widely the MGTs Zhdanov , Nojiri00 , Giacomozzi , Nashed ; however, one knows little about the self-gravitating topological structures [^1] Topological structures are nonlinear field configurations with stability due to topological properties. These objects arise particularly in field theory and particle physics. Generally speaking, one classifies these structures as kinks, vortices, and monopoles Vachaspati , Rajaraman , Manton , Vilenkin .. A significant class of topological structures are the topological vortices, which emerge from a complex scalar field theory coupled to an Abelian gauge field Vachaspati , Rajaraman , Manton , Vilenkin . Initially, Higgs used a massive scalar field in the mechanism that would later bear his name Higgs . Subsequently, the investigation of models coupled to an Abelian gauge field was introduced by Nielsen and Olesen Nielsen , adopting a flat spacetime. Generally speaking, vortex models preserve the Higgs mechanism and gauge invariance and are particularly relevant once these models are relativistic generalizations of the phenomenological Ginzburg-Landau theory for superconductivity James . Thus, the following question naturally arises: would it be possible for a BH-vortex system to emerge in an MGT? We aim to address this question in the development of this letter.
In this letter, we will explore the BH-vortex system within the $f(R)$ gravity. We begin by outlining the general framework for generating the BH-vortex system. Consequently, one investigates the BH and Hawking radiation in $f(R)$ gravity. Finally, the vortices surrounding the BH are analyzed, and our findings are announced.
II The BH-vortex system in modified $f(R)$ gravity
Let us explore the BH and magnetic vortex solutions in modified $f(R)$ gravity. This study is particularly relevant and helps us understand the interaction between the magnetic vortex and the BH. Although this research branch is still relatively unexplored, this study provides us valuable insights into the physical aspects of the BH-vortex system in MGT Bazeia1 , Bazeia2 . This approach also allows us to understand how MGTs, defined as functions of the Ricci scalar ($R$), influence gravitating matter. Furthermore, the Einstein-Hilbert theory corrected by quantum metric fluctuations leads us to consider the Maxwell-Higgs action in $f(R)$ gravity, viz.,
$S=\int\,d^{3}x\,\sqrt{-g}\bigg[-\frac{1}{16\pi}[R+f(R)]+\mathcal{L}_{\mathrm{mat}}\bigg],$ (1)
@@ II The BH-vortex system in modified $f(R)$ gravity
$\mathcal{L}_{\mathrm{mat}}=\frac{1}{2}(D_{\mu}\phi)^{\dagger}D^{\mu}\phi-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{\lambda}{4}(|\phi|^{2}-\nu^{2})^{2}.$ (2)
Here, $R$ is the Ricci scalar, $\phi$ a complex scalar field, and $F_{\mu\nu}$ the electromagnetic field tensor, defined as $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$, where $A_{\mu}$ is the gauge field. The parameter $\nu$ denotes the Vacuum Expectation Value (VEV). Furthermore, the notation $D_{\mu}\phi$ is the covariant derivative, which minimally couples the matter field $\phi$ to the gauge field $A_{\mu}$. One defines the covariant derivative as
@@
To obtain static and topological structures in three-dimensional spacetime Edery , let us adopt the metric
@@
where $h(r)$ is the metric function, $r$ and $\theta$ are, respectively, the radial and angular variables [^3] For more details on this metric, see Refs. Edery . In Refs. Edery , the authors used this metric to examine the vortex solution in Einstein’s gravity..
@@
$T_{\mu\nu}=$ $R_{\mu\nu}(1+f_{R})-\frac{1}{2}g_{\mu\nu}[R+f(R)]+(\nabla_{\mu}\nabla_{\nu}-g_{\mu\nu}\nabla_{\beta}\nabla^{\beta})f_{R},$ (10)
@@
To determine the profile of metric functions in gravity $f(R)$ theory, we will focus our study on cases where the effects of the presence of vortices are minimally perceptible. In other words, let us analyze regions outside the BH-Vortice system, i.e., sectors where the effects are negligible. In this regime, one can assume $T_{\mu\mu}=0$ [Eq. (11)]. This simplification allows us to obtain using Eq. (10), three independent differential equations, corresponding to the components $tt$, rrrr, and $\theta\theta$, i.e.,
@@
$2rh^{\prime}(r)f^{\prime}_{R}+2h^{\prime}(r)f_{R}-rh^{\prime\prime}(r)+rf(R)=0.$ (15)
@@
$R=-\frac{2h^{\prime}(r)}{r}-h^{\prime\prime}(r)$ (16)
and
$f(R)=-\alpha\left[\frac{2h^{\prime}(r)}{r}+h^{\prime\prime}(r)\right]$ (17)
@@
$(1+\alpha)rh^{\prime\prime}(r)=0\hskip 14.22636pt\text{with}\hskip 14.22636pt\alpha\neq-1.$ (18)
@@
$h(r)=r\mp r_{0}.$ (19)
Our purpose is to obtain BH-like compact objects. In pursuit of this, we consider only physically admissible solutions [Eq. (19)] from the theory, i.e., $h(r)=r-r_{0}$ [^6] The parameter $r_{0}$ may take arbitrary values. However, it is convenient to keep $r_{0}$ small once it is a comparison of the vortex matter and the horizontal event. That convenience is because minimal variations in $r_{0}$ induce abrupt changes in the magnetic vortices.. Indeed, the function $h(r)=r-r_{0}$ represents a simple and linear form of a curvilinear coordinate function in three-dimensional spacetime. In this framework, one can interpret the parameter $r_{0}$ as the BH mass. Thus, this function constitutes a modification of the radial term of the Schwarzschild-like metric. By analogy with the Schwarzschild-like BH, this metric function profile implies that the temporal term assumes negative values for $r<r_{0}$, suggesting the existence of an event horizon at $r=r_{0}$. Furthermore, in the limit $r\to r_{0}$, the temporal contribution vanishes, ensuring an essential feature of an event horizon Carroll  [^7] It is worth emphasizing that, in the limit $\alpha\to 0$, the action (1) reduces to the standard (naïve) formulation of the theory, viz., $S=\int\,d^{3}x\,\sqrt{-g}\left[-\frac{1}{16\pi}R+\mathcal{L}_{\text{matter}}\right]$. In this regime, the Einstein field equations of general relativity are recovered, i.e., $R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R=T_{\mu\nu}$, which admit recovering the Schwarzschild solution as a particular case. Alternatively, by parametrizing the theory through the inclusion of a $f(R)$-term arising from quantum fluctuations of the metric, viz., $f(R)=\alpha R$, one obtains a Schwarzschild-like theory, in which we reorganize all relevant parameters in the constant $r_{0}$. Furthermore, it is noteworthy that Eq. (15) reduces to $h^{\prime\prime}(r)=0$ in the limit of the standard theory, precisely matching the result found in the classical Schwarzschild solution. For further details, see Refs. Carroll ..
@@
Figure 1: Quadratic invariant of the Ricci tensor (blue curve) and the Kretschmann scalar (red curve) versus radial distance. This plot is valid for any value of $\alpha\neq-1$.
@@
When considering a linear function $f(R)$, one notes that the contribution arising from the quantum correction of the metric modifies the geometric sector, enabling adjustments to the effective Newtonian constant. This modification of the gravitational constant intensifies the deformations in the matter sector, altering the vortex’s matter configuration of the BH-vortex system. Consequently, this modification induces changes in the radiation emitted by the system as part of the matter collapses into the black hole. Therefore, this process affects phenomena on local scales, i.e., near the BH-vortex system  [^8] One can note this deformed behavior in Figs. 2(a) and 2(b). While in Einstein’s classical theory, the vortices exhibit profiles resembling kinks, in the $f(R)$ gravity theory announced in this work, deformations occur in the matter sector, specifically in the complex scalar and gauge fields. For further details, see Refs. Edery . .
II.0.1 On the Bekenstein–Hawking temperature